OpenStudy (anonymous):

What is the height of a trapezoid with one base of 8, another base of 14, and an area of 77?

OpenStudy (anonymous):

(8+14)(h)/2 = 77 (8+14)(h) =154 (22)(h)=154 h=7

OpenStudy (anonymous):

I LOVE YOU!

OpenStudy (anonymous):

I JUST GOT A '' A '' IN GEOMETRY!

OpenStudy (anonymous):

im your fan!

OpenStudy (anonymous):

Thanks